OpenStudy (babyslapmafro):

Please help me factor the following equation x^3+x^2-2=0

OpenStudy (babyslapmafro):

\[x^3+x^2-2=0\]

OpenStudy (ranga):

One way to solve cubic and higher order equations is to try out a few small values of x such as: 1, -1, 2, -2, etc. If I try x = 1 in x^3+x^2-2 it become 0 that means (x - 1) is a factor. Do synthetic division and figure out the quotient. It will be a quadratic expression which may be able to factor further.

OpenStudy (babyslapmafro):

ok i understand how (x-1) is a factor. but what do you mean by 'do synthetic division and figure out the quotient'?

OpenStudy (ranga):

(x - 1) is a factor means: The original expression x^3+x^2-2 = (x - 1){ some lower order expression } We have to figure out what the lower order expression is. So divide (x^3+x^2-2) by (x - 1) to figure out the lower order expression. This division you can do either the long division or if you are familiar with synthetic division you can use that method.

OpenStudy (babyslapmafro):

ok thanks

OpenStudy (ranga):

you are welcome.